OpenStudy (anonymous):

How many grams of potassium nitrate (KNO3) would form if 2.25 liters of a 1.50 molar lead nitrate Pb(NO3)2 solution reacts with 1.15 liters of a 2.75 molar potassium chromate K2CrO4 solution? Pb(NO3)2 + K2CrO4 PbCrO4 + 2KNO3 108 g 681 g 639 g 362 g @nechirwan

OpenStudy (nechirwan):

Balanced equation: Pb(NO3)2 + K2CrO4 → PbCrO4 + 2KNO3 1mol Pb(NO3)2 reacts with 1mol K2CrO4 to produce 2 mol KNO3 Check for limiting reactant Mol Pb(NO3)2 in 2.25L of 1.5M solution = 2.25*1.5 = 3.375 mol Mol K2CrO4 in 1.15L of 2.75M solution = 1.15*2.75 = 3.1625 mol The K2CrO4 is limiting From the equation: 1mol K2CrO4 will produce 2 mol KNO3 3.1625 mol K2CrO4 will produce 3.1625*2 = 6.325 mol KNO3 Molar mass KNO3 = 101.1g/mol 6.325 mol = 6.325*101.1 = 639 g